OpenStudy (yacoub1993):

What is the solution of the system? Use substitution. y = 6x + 2 3y - 18x = 12 A. No solution B. One Solution C. Two Solution D. Infinitely many solutions

OpenStudy (anonymous):

Screw substitution. Use elimination!

OpenStudy (anonymous):

y=6x+2 ...i 3y-18x=12....ii Now, substitute the value of y from equation i in equation ii then, 3(6x+2)-18x=12 18x+6-18x=12 6=12 which is false So no solution